# Data Readiness for AI: A 360-Degree Survey

### v1

@@ 4.1.15. FAIRness Score
FAIRness of a dataset for AI refers to the degree to which the dataset adheres to the principles of Findability, Accessibility, Interoperability, and Reusability. A FAIR dataset in the context of AI should be well-documented, easily discoverable, accessible, formatted in a way that facilitates integration and analysis and can be effectively reused for different AI applications.
@@
Example: Consider a FAIR dataset containing information about different types of cars for training an AI model tasked with predicting car prices. The FAIR principles applied to this dataset would resemble:
@@
  • Findability: The dataset should be easily discoverable with standardized metadata containing details about car attributes, including make, model, year, mileage, engine type, and other pertinent features. Clear information on the dataset’s source and data collection methodologies is also essential.
@@
  • Accessibility: The dataset’s accessibility is crucial, requiring hosting on a publicly available platform or a platform with controlled access, considering privacy or licensing constraints. Secure access should be ensured through proper authentication and authorization mechanisms.
@@
  • Interoperability: Ensuring seamless integration with AI systems and tools, the dataset should be structured and formatted according to established standards like CSV or JSON. A well-defined schema must accompany the dataset, specifying the meaning and format of each attribute. This promotes consistency and compatibility across diverse AI models and applications.
@@
  • Reusability: The dataset should come with clear usage licenses or permissions that outline how the dataset can be used. Additionally, comprehensive documentation should accompany the dataset, providing details about data collection procedures, preprocessing steps (such as data cleaning or feature engineering), and any potential biases or limitations in the data.
@@
By following these FAIR principles, the structured dataset of car information becomes a valuable resource for AI researchers and practitioners, facilitating the development of accurate car price prediction models and promoting transparent and ethical AI practices.
@@
Metrics in Literature: In their work, Wilkinson et al. (Wilkinson et al., 2018) introduced a comprehensive FAIRness measurement framework, aligning with the four FAIR sub-principles. The framework includes 14 universal metrics, corresponding to specific sub-principles, covering aspects such as identifier schemes, metadata accessibility, findability, access protocols, metadata longevity, knowledge representation languages, linking, adherence to standards, and provenance. This flexible approach, applicable across scholarly domains, facilitates the objective assessment and improvement of FAIRness in various digital resources. In a similar framework introduced by Clarke et al. (Clarke et al., 2019), FAIR metrics and FAIR rubrics play a key role, allowing users to associate digital resources with existing metrics. The authors emphasized the manual or automated quantification of FAIR metrics and contextual assessments using FAIR rubrics, empowering users to evaluate and enhance data correctness in diverse projects. DataONE(Data Observation Network for Earth) (Jones and Slaughter, 2019) is a community-driven initiative that has adopted metrics to measure FAIRness of research data. Based on the FAIR criteria, the DataONE FAIR suite generates comprehensive assessment scores based on the metadata.
@@
Summary: The FAIRness score, in the context of AI, measures the extent to which a dataset adheres to the principles of Findability, Accessibility, Interoperability, and Reusability. Both Wilkinson et al. and Clarke et al. created frameworks to assess the FAIRness of digital resources, aligned with FAIR sub-principles.
DRAI Metrics for Structured Data
@@ 4.2. Unstructured Data
Unstructured data, including textual, image, and audio data, present unique challenges in evaluating and ensuring readiness for AI applications. This section provides a brief overview of the existing metrics and scoring mechanisms employed to assess the suitability of unstructured data for AI. While the evaluation techniques for structured data are well-established, we will highlight the relevant measuring techniques from structured data that can be applied to unstructured data. By examining these evaluation methods, we gain insights into the quality, relevance, and accuracy of unstructured data, enabling us to make informed decisions when using such data in AI models. This exploration will facilitate a deeper understanding of the evaluation process and contribute to effectively integrating unstructured data in AI applications.
@@ 4.2.1. Textual Data
Firstly, let us examine some important dimensions of data readiness on textual data and explore the existing metrics and scoring mechanisms outlined in the literature for each dimension.
@@ 4.2.1.1 Lexical Diversity
Lexical diversity is a crucial metric for assessing the readiness of textual data for AI applications. It measures the richness, variety, and complexity of the vocabulary used within the text. By quantifying the diversity of word usage, lexical diversity offers insights into the level of linguistic expression, domain coverage, and potential challenges in understanding and processing textual data. A higher lexical diversity score indicates a broader range of words and linguistic patterns, providing a solid foundation for AI to learn from the data.
@@
Example: Consider two datasets: A with a low lexical diversity score and B with a high lexical diversity score. Dataset A comprises repetitive and limited vocabulary, such as a chatbot dialogue focused on a specific topic. The low lexical diversity suggests a constrained range of language, potentially limiting the chatbot’s ability to respond effectively to diverse user inputs. In contrast, Dataset B consists of a collection of news articles from various domains, exhibiting a wide range of vocabulary and language styles. The high lexical diversity in Dataset B indicates a greater readiness for AI applications, as it offers a more comprehensive representation of language usage, enabling models to generalize across different topics and understand a broader array of inputs.
@@
Metrics in Literature: TTR (Type-Token Ratio) introduced by Templin (Templin, 1957) is a metric used to measure lexical diversity in a text. It calculates the ratio of unique word types (vocabulary size) to the total number of tokens (words or other linguistic units) in the text. TTR provides an estimate of the text’s richness and variety of vocabulary. A higher TTR indicates greater lexical diversity, suggesting a more comprehensive range of word usage in the text.
@@
McCarthy et al.’s (McCarthy and Jarvis, 2010) metrics, vocd-D and HD-D, focus on measuring lexical diversity in textual data. vocd-D uses type-token ratios (TTR) from randomly selected text samples to derive a D coefficient, representing lexical diversity. HD-D, on the other hand, employs the hypergeometric distribution to directly calculate the probabilities of word occurrence in randomly selected samples, resulting in the HD-D index. While both metrics assess lexical diversity, vocd-D uses random sampling and the D coefficient, whereas HD-D approximates results for all possible word arrangements. The correlation between HD-D and vocd-D is high, offering alternative perspectives on measuring lexical diversity. Additionally, McCarthy introduces MTLD (Measure of Textual Lexical Diversity) (McCarthy, 2005), which quantifies lexical diversity by considering unique words and segment length. MTLD is beneficial for assessing the lexical diversity of longer texts, providing insights into vocabulary richness throughout the text.
@@
Summary: Lexical diversity, necessary for assessing linguistic richness in textual data for AI, is measured by metrics like TTR and vocd-D, evaluating the ratio of unique word types to total tokens with standardized scores. HD-D offers a probability-based alternative for assessing diversity, while MTLD, dividing text into segments, calculates average segment length where vocabulary richness falls below a threshold.
@@ 4.2.1.2 Term Importance
Term importance is used to assess the readiness of textual data for AI applications by evaluating the significance of individual terms within the text. It measures terms’ relevance and impact in capturing the text’s essence and meaning. Term importance considers various factors, such as the frequency of a term within the text, its rarity across the entire dataset or corpus, and its discriminative power in distinguishing the text from others. By assigning weights or scores to terms based on their importance, this metric enables AI models to focus on key terms that carry valuable semantic information and discard less informative or common terms, assisting in feature selection, document ranking, or topic extraction.
Example: In a dataset of news articles, the term “pandemic” might be considered highly important due to its relevance in conveying crucial information about current events. On the other hand, common words like “the” or “and” would be assigned lower importance scores as they provide little discriminative power or unique information. AI models can prioritize and focus on the most significant terms by analyzing term importance, enabling better understanding, classification, or summarization of textual data.
Metrics in Literature: TF-IDF (Term Frequency-Inverse Document Frequency) is a widely used quantitative metric to evaluate the importance of words in a document or collection of documents (Ramos (Ramos et al., 2003), Simha (Simha, 2021), Qaiser et al. (Qaiser and Ali, 2018)). It combines term frequency (TF), measuring word occurrence in a document, and inverse document frequency (IDF), assessing word rarity across the entire corpus. TF-IDF quantifies a word’s significance by considering its frequency in a document and its discriminative power across the corpus. The TF-IDF score is computed by multiplying TF and IDF values for each word. High TF-IDF scores indicate words that are both frequent in a document and rare in the corpus, making them essential for the context. This metric aids in identifying essential features and characteristics, enabling information retrieval, document classification, and keyword extraction in natural language processing.
Summary: Term importance is a metric used to assess the significance of individual terms in textual data for AI applications. TF-IDF is a widely used metric that combines term frequency and inverse document frequency to quantify term importance.
@@ 4.2.1.3 Readability Score
The readability score is a quantitative metric used to assess textual data complexity and ease of understanding, enabling effective preparation for AI. It measures various linguistic factors, such as sentence length, word choice, and syntactic structure, to determine the readability of a text. By considering these factors, readability scores provide valuable insights into the suitability of text for different target audiences and applications. AI applications can be optimized by selecting appropriate training data using readability scores, ensuring that the content aligns with the desired level of comprehension and avoids potential barriers to understanding. This metric is important in enabling the development of more accessible and contextually appropriate language models.
@@
Example: Consider a scenario where an AI model is trained to generate educational content for elementary school students. In this case, readability scores can be used to assess the complexity of different texts and select appropriate training data. By analyzing factors such as sentence length, vocabulary difficulty, and grammatical complexity, the readability scores can help identify texts that align with the target audience’s reading abilities. This ensures that the AI model is trained on comprehensible and engaging content for young learners, promoting effective knowledge transfer and enhancing the overall learning experience. The AI model can be fine-tuned by leveraging readability scores to generate age-appropriate and easily understandable educational materials catering to elementary school students’ unique needs and capabilities.
@@
Metrics in Literature: The Flesch-Kincaid Grade Level, introduced by Rudolf Flesch (Flesch, 1986) calculates the approximate grade level needed to comprehend a given text based on the average number of words per sentence and the average number of syllables per word. The resulting score represents the education level required to understand the text, with lower values indicating easier readability. This metric provides a standardized measure for assessing text comprehension and enables content tailoring to suit specific audience reading abilities.
@@
The Coleman-Liau Index, developed by Coleman and Liau (Coleman and Liau, 1975), is another readability scoring method that assesses the reading level of a text based on factors like letter count and sentence length. Unlike the Flesch-Kincaid Grade Level, which considers syllable count, the Coleman-Liau Index calculates the grade level based on the average number of letters per 100 words and the average number of sentences per 100 words. A score of 5 on the index indicates that the text is at a reading level equivalent to that of a fifth grader in the US schooling system. It is widely used in schools and provides a quick measure of readability.
@@
Furthermore, the Gunning Fog Index introduced by Robert Gunning Associates (rea, nd) offers an additional perspective on the readability of textual data for AI model training. Unlike the Coleman-Liau Index and the Flesch-Kincaid Grade Level, which focus on sentence and letter count, the Gunning Fog Index considers both the percentage of complex words and the sentence length. It generates a score between 0 and 20, with lower scores representing easier readability.
@@
Summary: The readability score is a quantitative metric used to assess the complexity of textual data and ease of understanding. Popular readability metrics include the Flesch-Kincaid Grade Level, which calculates the grade level needed to comprehend a text based on average words per sentence and syllables per word. The Coleman-Liau Index assesses reading level based on average letter count and sentence length. The Gunning Fog Index evaluates readability by considering the percentage of complex words and sentence length.
@@ 4.2.1.4 Topic Coherence
This measure evaluates the readiness of textual data for AI by assessing the logical and semantic connectedness within a set of topics or a document. It quantifies the degree to which words within a topic exhibit meaningful relationships and contribute to a coherent theme. A higher coherence score indicates stronger semantic coherence, indicating that the words are closely related and provide a clearer understanding of the topic. By evaluating topic coherence, AI practitioners can ensure that the textual data is well-structured, coherent, and ready for AI model training, promoting more accurate and meaningful text generation and facilitating better comprehension and usage of the data by AI algorithms.
@@
Example: Consider a collection of news articles about technology trends. Topic coherence can be measured to evaluate the readiness of this textual data for AI. Data can be segmented into topics like “Artificial Intelligence,” “Blockchain Technology,” and “Internet of Things” by applying topic modeling techniques. Topic coherence analysis assesses the semantic relationships between words within each topic. A high coherence score would indicate that words within a topic, such as “machine learning,” “algorithm,” and “predictive analytics” in the “Artificial Intelligence” topic, are closely related and contribute to a coherent theme. This demonstrates that the textual data is well-prepared for AI, as it exhibits clear and meaningful topic structures, allowing the models to understand better and generate text related to specific technology trends.
@@
Metrics in Literature: The CV coherence score, proposed by R”oder et al. (Röder et al., 2015), is a metric for quantifying topic coherence in text by assessing the semantic similarity between words within a topic. It is computed using techniques like Latent Dirichlet Allocation (Blei et al., 2003) to extract topics from the text and then measures the pairwise word similarity within each topic to evaluate how well the words contribute to a coherent theme. Higher scores indicate stronger semantic relatedness and better topic coherence. Despite its popularity, the CV score has limitations, such as sensitivity to topic size, potential mismatch with human judgment, and inability to capture higher-level coherence aspects.
@@
The UMass coherence score, introduced by Mimno et al. (Mimno et al., 2011), is another metric used to assess the topic coherence of a set of topics extracted from a text corpus. It evaluates coherence by considering the probability of word co-occurrences within topics. The score is computed by summing the logarithm of the co-occurrence probabilities of all word pairs within and across topics. A higher UMass coherence score indicates stronger word co-occurrence patterns and better topic coherence, while a lower score suggests weaker word associations and less coherent topics. Compared to the CV coherence score, the UMass coherence score offers a more reliable evaluation of topic coherence, accounting for topic size, aligning better with human judgment, and directly measuring word co-occurrence probabilities.
@@
Like other coherence scores, the UCI coherence score introduced by Newman et al. (Newman et al., 2010) is a metric used to assess the coherence of topics generated by a topic model. It measures the semantic relatedness and meaningful connections between words within a topic by calculating the semantic association between word pairs based on their co-occurrence in sliding windows. The score is computed using a specific equation considering the probabilities of observing individual words and word pairs in the sliding window. Higher UCI coherence scores indicate stronger associations between word pairs and better topic coherence.
@@
Summary: Topic coherence is a metric used to evaluate the logical and semantic connectedness within a set of topics or a document. Coherence scores, such as the CV coherence score, UMass coherence score, and UCI coherence score, quantify the semantic similarity or word co-occurrence probabilities within topics to assess their coherence. Higher coherence scores indicate stronger semantic relatedness and better topic coherence.

### v2

@@ 3.1.15. FAIR Principle Compliance Score
FAIR compliance of a dataset for AI refers to the degree to which a dataset adheres to Findability, Accessibility, Interoperability, and Reusability (FAIR) principles (FAI, 2024). In AI, a dataset should be well-documented, easily discoverable, accessible, formatted in a way that facilitates integration and analysis, and can be effectively reused for AI applications.
@@
Example: Consider a dataset containing information about different types of cars for training an AI model tasked with predicting car prices. The FAIR principles applied to this dataset would be:
@@
  • Findability: The dataset should contain a unique identifier to be easily discoverable with standardized metadata containing details about car attributes, including make, model, year, mileage, engine type, and other existing features. Clear information on the dataset’s source and data collection methodologies is also essential.
@@
  • Accessibility: The dataset’s accessibility requires a platform with controlled access, considering privacy or licensing constraints. Secure access should be given through proper authentication and authorization methods.
@@
  • Interoperability: The dataset should be structured and formatted according to established storage standards like CSV or JSON to ensure easier integration with AI systems and tools. A well-defined schema must be included with the dataset, specifying the meaning and format of each attribute. This promotes consistency and compatibility across diverse AI models and applications.
@@
  • Reusability: The dataset should come with clear usage licenses or permissions that outline how the dataset can be used. Additionally, comprehensive documentation should be included with the dataset, providing details about data collection procedures, preprocessing steps (such as data cleaning or feature engineering), and any potential biases or limitations in the data.
@@
By following these FAIR principles, a structured dataset of car information becomes a valuable resource for AI researchers and practitioners. It facilitates the development of accurate car price prediction models and promotes transparent and ethical AI practices.
@@
Metrics in Literature: Wilkinson et al. (2018) introduced a comprehensive FAIR compliance measurement framework, aligning with the four FAIR sub-principles. The framework includes $14$ universal metrics corresponding to specific sub-principles, covering aspects such as identifier schemes, metadata accessibility, findability, access protocols, metadata longevity, knowledge representation languages, linking, adherence to standards, and provenance. This flexible approach facilitates the objective assessment and improvement of FAIR compliance in various digital resources applicable across scholarly domains. In a similar framework introduced by Clarke et al. (2019), FAIR metrics and FAIR rubrics play an important role, allowing users to associate digital resources with existing metrics. The authors emphasized the manual or automated quantification of FAIR metrics and contextual assessments using FAIR rubrics. This empowers users to evaluate and enhance data correctness in diverse projects. DataONE (Data Observation Network for Earth) (Jones and Slaughter, 2019) is a community-driven initiative that has adopted metrics to measure FAIR principle compliance of research data. Based on the FAIR criteria, the DataONE FAIR suite generates comprehensive assessment scores based on the metadata.
Impact on AI: A dataset’s compliance with the FAIR principles can impact AI by enhancing data management and accessibility. By ensuring data is easily discoverable and accessible, AI systems can be trained on diverse and comprehensive datasets, leading to robust and accurate models. Interoperability allows different AI systems to work together seamlessly, while reusability enables researchers to build upon existing datasets and models, enabling reproducible innovation and efficiency. Moreover, FAIR compliance promotes collaboration and transparency that accelerates AI research and ensures the reproducibility of results. It is essential for building trust in AI technologies.
@@
Summary: The FAIR score measures the extent to which a dataset adheres to the principles of Findability, Accessibility, Interoperability, and Reusability. Both Wilkinson et al. and Clarke et al. created frameworks to assess the FAIR compliance of digital resources aligned with FAIR sub-principles. Several other FAIR compliance score evaluation websites use the same $14$ principles that Wilkinson et al. defined.
Data Readiness Metrics for Structured Data In evaluating structured data readiness for AI applications, we considered metrics spanning completeness, outliers, mislabels, duplicates, feature relevancy, class imbalance, class separability, discrimination index, data split ratio, data point impact, correctness, timeliness, privacy leakage, sample size, and FAIR compliance score.
@@ 3.2. Unstructured Data
Unstructured data, including textual, image, and audio data, present unique challenges in evaluating and ensuring readiness for AI applications. In this section, we provide a brief overview of the metrics and scoring mechanisms (shown in the right half of Table 2) used to assess the suitability of unstructured data for AI. While the evaluation techniques for structured data are well-established, we will highlight the relevant measuring techniques from structured data that can be applied to unstructured data. By examining these evaluation methods, we gain insights into the readiness, relevance, and accuracy of unstructured data. It enables us to make informed decisions when using such data in AI models.
@@ 3.2.1.1 Lexical Diversity
Lexical diversity measures the richness, variety, and complexity of the vocabulary used within the text. It offers insights into the level of linguistic expression, domain coverage, and potential challenges in understanding and processing textual data. A higher lexical diversity score indicates a broader range of words and linguistic patterns, providing a solid foundation for AI to learn from the data.
@@
Example: Consider two datasets: A with a low lexical diversity score and B with a high lexical diversity score. Dataset A has a repetitive and limited vocabulary, such as a chatbot dialogue focused on a specific topic. The low lexical diversity suggests a constrained range of language, potentially limiting the chatbot’s ability to respond effectively to diverse user inputs. In contrast, dataset B consists of a collection of news articles from various domains, exhibiting a wide range of vocabulary and language styles. The high lexical diversity in dataset B indicates a greater readiness for AI applications. It offers a more comprehensive representation of language usage, enabling models to generalize across different topics and understand a broader range of inputs.
@@
Metrics in Literature: Type-Token Ratio (TTR) introduced by Templin (1957) measures lexical diversity in textual data. It calculates the ratio of unique word types (vocabulary size) to the total number of tokens (words or other linguistic units) in the text. TTR provides an estimate of the text’s richness and variety of vocabulary. A higher TTR indicates greater lexical diversity, suggesting a more comprehensive range of word usage in the text.
@@
McCarthy et al.’s (McCarthy and Jarvis, 2010) metrics, vocd-D and HD-D, focus on measuring lexical diversity in textual data. vocd-D uses type-token ratios (TTR) from randomly selected text samples to derive a D coefficient to represent lexical diversity. HD-D, on the other hand, employs the hypergeometric distribution to directly calculate the probabilities of word occurrence in randomly selected samples, resulting in the HD-D index. While both metrics assess lexical diversity, vocd-D uses random sampling and the D coefficient, whereas HD-D approximates results for all possible word arrangements. The correlation between HD-D and vocd-D is high, offering alternative methods of measuring lexical diversity. Additionally, McCarthy introduces the Measure of Textual Lexical Diversity (MTLD) (McCarthy, 2005), which quantifies lexical diversity by considering unique words and segment length. MTLD assesses the lexical diversity of longer texts, providing insights into vocabulary richness throughout the text.
Impact on AI: Measuring lexical diversity can significantly impact AI by influencing how language models are developed and evaluated. In AI, particularly in Natural Language Processing (NLP) and machine translation, assessing lexical diversity can help identify biases and limitations in language models. Reviriego et al. (2023) state that AI-generated texts often exhibit lower lexical diversity compared to human-generated texts, leading to a feedback loop where AI models trained on such data become less effective over time. Ensuring high lexical diversity in AI outputs can improve the quality and accuracy of translations and other language-based AI applications. This ultimately contributes to more inclusive and representative AI systems.
@@
Summary: Lexical diversity, necessary for assessing linguistic richness in textual data, is measured by metrics like TTR and vocd-D, evaluating the ratio of unique word types to total tokens with standardized scores. HD-D offers a probability-based alternative for assessing diversity. MTLD divides the text into segments and calculates the average segment length where vocabulary richness falls below a threshold.
@@ 3.2.1.2 Term Importance
Term importance evaluates the significance of individual terms in textual data. It measures the relevance and impact of terms in capturing the essence and meaning. Term importance considers various factors, such as the frequency of a term within the text, its rarity across the entire dataset or corpus, and its discriminative power in distinguishing the text from others. By assigning weights or scores to terms based on their importance, this metric enables AI models to focus on key terms that carry valuable semantic information and discard less informative or common terms, assisting in feature selection, document ranking, or topic extraction.
Example: In a dataset of news articles, the term “pandemic” might be considered highly important due to its relevance in conveying crucial information regarding COVID-19. On the other hand, common words like “the” or “and” would be assigned lower importance scores as they provide little discriminative power or unique information. AI models can prioritize and focus on the most significant terms by analyzing term importance, enabling better understanding, classification, or summarization of textual data.
Metrics in Literature: TF-IDF (Term Frequency-Inverse Document Frequency) is a widely used quantitative metric to evaluate the importance of words in a document or collection of documents (Ramos et al. (2003), Simha (2021), Qaiser and Ali (2018)). It combines term frequency (TF), measuring word occurrence in a document, and inverse document frequency (IDF), assessing the rarity of a word’s appearance across the entire corpus. TF-IDF quantifies a word’s significance by considering its frequency in a document and its discriminative power across the corpus. The TF-IDF score is computed by multiplying TF and IDF values for each word. High TF-IDF scores indicate that words that are both frequent in a document and rarely appear in the corpus, which makes them essential for the context. This metric helps identify essential features and characteristics, enabling information retrieval, document classification, and keyword extraction in natural language processing.
Impact on AI: Measuring term importance is needed for tasks such as information retrieval, text summarization, and sentiment analysis, where distinguishing between relevant and irrelevant terms can greatly impact the accuracy and relevance of the output. By measuring term importance accurately, AI systems can prioritize critical information that leads to precise and contextually relevant results. Furthermore, TF-IDF aids in reducing computational resource requirements in AI training by focusing on significant terms.
Summary: Term importance assesses the significance of individual terms in textual data for AI applications. TF-IDF is a widely used metric that combines term frequency and inverse document frequency.
@@ 3.2.1.3 Readability Score
Readability score is a quantitative metric used to assess textual data complexity and ease of understanding, enabling effective preparation for AI. It measures various linguistic factors, such as the length of sentences, choice of words, and syntactic structure, to determine the readability of a text. By considering these factors, readability scores provide valuable insights into the suitability of text for different target audiences and applications. AI applications can be optimized by selecting appropriate training data using readability scores, ensuring that the content aligns with the desired level of comprehension and avoids potential barriers to understanding. This metric is important in enabling the development of more accessible and contextually appropriate language models.
@@
Example: Consider a scenario where an AI model is trained to generate educational content for an elementary school. In this case, readability scores can be used to assess the complexity of different texts and select appropriate training data. The readability scores can help identify texts that align with the target audience’s reading abilities by analyzing factors such as sentence length, vocabulary difficulty, and grammatical complexity. This ensures that an AI model is trained on comprehensible and engaging content for young learners, promoting effective knowledge transfer and enhancing the overall learning experience.
@@
Metrics in Literature: The Flesch-Kincaid Grade Level, introduced by Flesch (1986) estimates the approximate education (grade) level needed to comprehend a given text using the average number of words in sentences and that of syllables in words. The resulting score represents the education level required to understand the text. Lower grade values indicate higher/easier readability. This metric provides a standardized measure for assessing text comprehension and enables content tailoring to suit specific audience reading abilities.
@@
The Coleman-Liau Index, developed by Coleman and Liau (1975), is another readability scoring method that assesses the reading level of a text based on factors like letter count and sentence length. Unlike the Flesch-Kincaid Grade Level, which considers syllable count, the Coleman-Liau Index calculates the grade level based on the average number of letters and sentences per $100$ words. A score of $5$ on the index indicates that the text is at a reading level equivalent to that of a fifth grader in the US schooling system. It is widely used in schools and provides a quick measure of readability.
@@
Furthermore, the Gunning Fog Index introduced by Robert Gunning Associates (rea, nd) offers an additional perspective on the readability of textual data for AI model training. Unlike the Coleman-Liau Index and the Flesch-Kincaid Grade Level, which focus on sentence and letter count, the Gunning Fog Index considers both the percentage of complex words and the sentence length. It generates a score between 00 and $20$, with lower scores representing easier readability.
Impact on AI: By evaluating readability, developers can adjust the complexity of the training data to match the intended audience’s comprehension level, which is particularly important in fields like education and healthcare, where clear communication is essential. By ensuring that training data is readable, AI models can avoid perpetuating biases that arise from overly complex or inaccessible language. This will also reduce disparities in information access.
@@
Summary: The readability score is a quantitative metric used to assess the complexity of textual data and ease of understanding. Popular readability metrics calculate the grade level needed to comprehend a text based on average words per sentence and syllables per word and by considering the percentage of complex words and sentence length.
@@ 3.2.1.4 Topic Coherence
This measure evaluates the readiness of textual data for AI by assessing the logical and semantic connectedness within a set of topics or a document. It quantifies the degree to which words within a topic exhibit meaningful relationships and contribute to a coherent theme. A higher coherence score shows stronger semantic coherence, indicating that the words are closely related and provide a clearer understanding of the topic. By evaluating topic coherence, AI practitioners can ensure that the textual data is well-structured, coherent, and ready for AI model training. This would promote accurate and meaningful text generation and facilitate better comprehension and usage of data by AI algorithms.
@@
Example: Consider a collection of news articles about technology trends. Topic coherence can be measured to evaluate the readiness of this textual data for AI. Data can be segmented into topics like “Artificial Intelligence,” “Blockchain Technology,” and “Internet of Things” by applying topic modeling techniques. Topic coherence analysis assesses the semantic relationships between words within each topic. A high coherence score would indicate that words within a topic, such as “machine learning,” “algorithm,” and “predictive analytics” in the “Artificial Intelligence” topic, are closely related and contribute to a coherent theme. This demonstrates that the textual data is well-prepared for AI, as it exhibits clear and meaningful topic structures.
@@
Metrics in Literature: R”oder et al. propose the “CV coherence score” (Röder et al., 2015) to quantify topic coherence in textual data by assessing the semantic similarity between words within a topic. This is computed using Latent Dirichlet Allocation (Blei et al., 2003) to extract topics from the text and then measure the pairwise word similarity within each topic to evaluate how well the words contribute to a coherent theme. Higher scores indicate stronger semantic relatedness and better topic coherence. Despite its popularity, the CV coherence score has limitations, such as sensitivity to topic size, potential mismatch with human judgment, and inability to capture higher-level coherence aspects.
@@
Mimno et al. introduce “UMass coherence score” (Mimno et al., 2011) to assess the topic coherence of a set of topics extracted from a text corpus. It evaluates coherence by considering the probability of word co-occurrences within topics. The score is computed by aggregating the logarithm of the co-occurrence probabilities of all word pairs within and across topics. A higher UMass coherence score indicates stronger word co-occurrence patterns and better topic coherence, while a lower score suggests weaker word associations and less coherent topics. Compared to the CV coherence score, the UMass coherence score offers a more reliable evaluation of topic coherence, accounting for topic size, aligning better with human judgment, and directly measuring word co-occurrence probabilities.
@@
Newman et al. proposed “UCI coherence score” (Newman et al., 2010) to assess the coherence of topics generated by a topic model. This measures the semantic relatedness and meaningful connections between words within a topic by calculating the semantic association between word pairs based on their co-occurrence in sliding windows. The score is computed using a specific equation considering the probabilities of observing individual words and word pairs in a sliding window. Higher UCI coherence scores indicate stronger associations between word pairs and better topic coherence.
Impact on AI: By evaluating and optimizing for topic coherence, developers can ensure that the AI models produce topics that align better with human understanding, which will improve the semantic interpretability of the model outputs. This evaluation can lead to accurate and reliable AI systems, as coherent topics are more likely to represent genuine patterns in the data rather than random associations.
@@
Summary: Topic coherence is a metric used to evaluate the logical and semantic connectedness within a set of topics or a document. Coherence scores, such as the CV, UMass, and UCI coherence scores, quantify the semantic similarity or word co-occurrence probabilities within topics to assess their coherence.